Debby and Cindy are sisters who have decided they want to lose weight for their upcoming family wedding. Debby was a healthy weight in high school and was very active in sports. She has gained weight over the past few years since graduating from college 5 years ago. She is currently 5'6" and weighs 160 lb. Cindy was overweight as a child and has continued to gain more weight as an adult. She is 5'4" and weighs 230 lb. Debby decides she is going to go on a High Protein/ Low Carbohydrate diet to attempt to lose 25 lbs in 2 months to get back to her high school weight. Cindy has tried diets in the past and has found that she always gains the weight back when she goes off the diet, so she decides to consult her doctor for more aggressive treatment. She discovers that she has high blood sugar, high blood pressure and a family history of heart disease. The doctor discusses the options of medications as well as gastric surgery with her. Cindy has heard about Phentermine from some family members, but she found the doctor will not prescribe that for her but he did discuss the option of gastric surgery with her.

1. Calculate and show Debby and Cindy's BMI's. Do they both need to lose weight? If they wanted to lose weight, what is a realistic amount of weight for each of them to lose in 2 months.
Answer #1

BMI calculation

Multiply weight in pounds by 705 then divide the result by height in inches , then divide again by height in inches . The result is BMI.

1. * BMI of Debby is

160× 705= 112800

112800÷ 66 =1709.09

1709.09 ÷ 66= 25.8

25.8 is debby's bmi . It indicates she  is overweight.

BMI between 25 -29.9 is considered as overweight.

* BMI of cindy

230 ×705 =162150

162150 ÷ 64 = 2533.59

2533.59 ÷ 64 =39.5

39.5 is cindy's BMI. It indicates she is obese.

BMI between 30 and above is considered as obese.

Both of them need to lose weight.

Normal BMI is from 18.5 to 24.9.

To become 24.9 range debby has to lose 6 pounds and cindy has to lose 85 pounds.

Lose 1 - 2 lbs per week ie, 4- 8 lbs per month is healthy goal.

So debby can lose her weight in a month.

Based on this healthy goal cindy can lose only 16 pounds  in 2 months .

If cindy undergo surgery then she might lose 20 pounds per month shortly after the surgery and 10 pounds per month and then 5 pounds per month.